# Improved Tangential Interpolation-based Multi-input Multi-output Modal Analysis of a Full Aircraft

### v1

@@ 3.2 Computational performance
As already shown in [10] for the SIMO modal parameters identification, N4SID is orders of magnitude slower than LF and LSCE. LSCE is found to be the fastest method, but it is clear that it does not serve a purpose as it cannot identify modes. However, it’s important to see that the iLF is an order of magnitude faster than the traditional implementation of LF. In addition, the conventional LF shows a slight dependency on model order, which is not the case for the iLF. The regressed line equation verifies this in Figure 3.
@@ 3.3 Investigation of noise effects
After having preliminary validated the iLF identification against the traditional LF and the benchmark methods, the next step is to assess its robustness to noise. To achieve this, the time series are corrupted with different levels, with respect to the standard deviation of the signal, of Additive white Gaussian noise (AWGN). The selected levels are 0.1, 0.5, 1, 1.5 and 2% (note that both the input and output signals are corrupted with noise). Then, the corrupted time series are used to obtain the FRFs, which in turn are fed to iLF to obtain the modal parameters. The validation process is carried out in two steps: (i) the minimum order = 16 is considered and (ii) the stable modes between orders 32 and 60 are examined.
@@
From Figure 4 it is clear that the first 15 modes are identified within acceptable error margins. For example, the maximum error in terms of is less than 0.5%, less than 8.5% for and the lowest MAC value of is 0.74, being 0.8 the minimum for an acceptable correlation. In addition, the first nine modes (exception made for mode #5 for the 1.5% noise level) are coherent with the analytically derived values for all noise levels. As expected, the maximum discrepancies are found for the higher modes. This is a common trend for LF-derived techniques and it was also found in the SIMO application of LF reported in [11]. The most significant issues in modal identification precision are found for the higher modes and the modes close in frequency.
The results of the identification of stable modes from the noised signals are shown in Figure 5. Here the results are similar to that of the noiseless case, with minimal variations (less than 0.06%) in and virtually no changes in the identified (differences are all of 0%) and (all MAC values are 1). In conclusion, the computational improvement of the iLF with respect to the conventional LF does not hamper its precision and robustness to noise. In particular, it offers the same precision as industry-standard methods, like N4SID, on the numerical cases analysed. In addition, it can identify systems which standard methods, like LSCE, struggle with, as it was for the case studied. Further, experimental dataset verification and validation are carried out in the remainder of the article.
@@ 4 The BAE Systems Hawk T1A aircraft benchmark
As of the writing of this article, the experimental data from the BAE Systems Hawk T1AHawk T1A aircraft has been recently published by researchers from the Dynamics Research Group at the University of Sheffield.
The full aircraft, a Hawk T1A formerly in service with the British Royal Air Force and now decommissioned, is depicted in Figure 6. The craft is shown at the University of Sheffield’s Structural Dynamics Laboratory for Verification and Validation (LVV) facilities (Sheffield, UK), where it was tested.
Importantly, the data used in this article are firstly described in two separate documents: [25] and [26]. They focus on, respectively, the starboard wing and the whole Hawk airframe structure. Both were used here for experimental validation.
These two related datasets serve as a compelling benchmark for dynamic System Identification techniques. Nevertheless, as discussed in this Section, some precautions have been considered to ensure the full comparability of data from the two experimental campaigns.
@@ 4.1 Wing experimental setup and data description
A total of 54 PCB Piezotronics accelerometers (nominal sensitivity: 10 mVg-1) are mounted on the lower and upper surface of the starboard wing, as shown in Figure 7. They are arranged in a grid of three and two lines, respectively, running parallel to the leading edge. Extra sensors are placed at the root boundary condition locations and by the undercarriage door. One Tira TV 51140-MOSP modal shaker, directly attached to a PCB Piezotronics 208C02 force transducer with a sensitivity of 11.11 mVN-1, applies the input driving force on the wing lower surface. Specifically, the acquisitions corresponding to burst random excitation with 0.4 V input voltage are considered in this study.
As can be inferred, this dataset only includes one input source. Therefore, it was used to obtain the single input single output readings to compare the SISO and SIMO capabilities of the iLF while getting baseline values for system modes.
4.1.1 Whole airframe setup and data description
@@
Regarding the input signal and amplitude, for comparability with the wing dataset, the white Gaussian noise data are used, with the same 0.4 V LMS input voltage. In this case, five modal shakers are attached to the starboard wing (i.e. the one already discussed above), the port wing, plus the starboard horizontal stabiliser, port horizontal stabiliser, and vertical stabiliser. The exact locations are marked in orange in Figure 8.
@@
The two sides of the horizontal stabiliser are populated with two rows of five uniaxial sensors at both the rear and front side, totalling 20 per side (located exclusively on the upper; see Figure 8(c));
@@
Similarly, the vertical stabiliser sensors are placed only on the starboard side, again with 10 sensors equally distributed along the leading and trailing edge (Figure 8(d));
@@
Another set of four high-sensitivity (100 mVg-1) uniaxial PCB Piezotronics accelerometers, are placed on the fuselage midpoint on the upper, lower, port, and starboard sides (F_MLU, F_MLL, F_MLP, and F_MLS in Figure Figure 9(a));
@@
The 86th sensor, a triaxial accelerometer laid on the ground as a reference, is obviously not used here for SI. Apart from this, all the other output channels are selected for MIMO System Identification, in conjunction with the five input channels. Conversely, the fibre-Bragg grating strain gauges - described in [28] - were not considered here.
@@ 4.1.2 Pre-processed FRFs
estimators are indeed appreciated for their capability to reduce output measurement noise, which more significantly affects the valleys of FRFs with signal leakage. However, it is well-known that, while the FRF estimators allow for averaging over multiple acquisitions and provide more accurate estimates in absolute value, this is achieved at the cost of losing the phase information (see e.g. [29] Chapter 3.13). The LF and its variant iLF are intended for fitting over complex-valued Transfer Functions, as will be shown for the numerical example and the first set of experimental data; however, they can be applied in this case of real-valued functions, as it is shown in the results of the second dataset.
@@ 4.1.3 Further considerations
Before continuing, there are two key aspects worth mentioning.
The first one is that the sensor layout for the starboard wing used in [28] is similar but not identical to the one of [27], as e.g. only two lines of sensors were used on the upper side (one on the lower side). As it is better discussed in the Results Section, the nearest output channels were selected in order to compare SISO, MISO, and MIMO tests between these two datasets.
The second aspect is that for the experiments on the whole airframe with WGN input, the excitation bandwidth was 5–256 Hz and the sampling rate was 2048Hz. On the other hand, the first experimental campaign used a sampling frequency of 512 Hz but it was bandlimited between 5-256 Hz. Therefore, all analyses have been here limited to Hz.
@@ 4.2 Modal parameters estimation
Previous MIMO modal testing work on the aircraft does not explicitly report any results to be used as an external benchmark [26]. Nevertheless, data is available from the SIMO modal testing of the aircraft starboard wing [25]. However, the results presented here are only for the SISO identification via RFP from the FRF for channel LTC-05 (Figure 10(a)), with no further numerical values included. It is clear that, for getting an idea of where (in frequency) the modes of the full aircraft are, this approach is limited by the data considered. In addition, it should be noted that the second to last mode (mode #22 according to [25]) is very likely a repetition of the previous mode (potentially a typo). Nevertheless, it is possible to retrieve its actual value, as the authors shared the code used to obtain those results [30].
In [25], ranges for the natural frequencies of the starboard wing are supplied to be used to identify the modes via RFP; hence, the identification is carried out in a peak-picking fashion rather than using higher-order models. Thus, the idea is to compare this SISO result with a MISO result from the same, or close, sensor. To do this, the sensors position graphs of the starboard wing supplied in [25] (SIMO) and [26] (MIMO) are superimposed in Figure 10(a). There, it is shown that the shaker is located in, almost, the same position in the two tests; however, the sensor row of LTC-05 is not present in the MIMO sensor layout. Hence, another sensor, with a counterpart in the MIMO test, is identified. In the SIMO test, this sensor is known as LLC-07, while as SW_LC1 in the MIMO test. Please note that in all cases considered the shaker input signal is random white noise at 0.4 V.
@@
The expectation is that N4SID and iLF will not perform as well as RFP in this task as they are more holistic methods implemented for more typical scenarios with multiple outputs and inputs, rather than being used in a peak-picking fashion. Then, the SIMO and MISO analyses outlined above follow.
The identified and , for the different cases and sensors, are shown, respectively, in Tables 2 and 3. The results marked with (*) are those erroneously reported in [25] and retrieved with the authors-provided [30] RFP implementation. Please note that only frequencies [5-160] Hz are inspected in all cases for coherence with [25].
@@
The results presented in [25] for the band-limited SISO identification via RFP of the LTC-05 sensor of the starboard wing dataset show that 23 modes are found for the wing. Of course, these modes are representative of a local state and more modes are found when the full aircraft is tested. After all the SISO, SIMO, MISO, and MIMO identifications 44 modes are identified in at least one of the cases. Hence, from now on, these modes are referred to with the numbering shown in Tables 2 and 3.
In terms of for the SISO identifications, N4SID and iLF always match the frequencies identified by RFP in LTC-05 and LLC-07, except for modes # 20, 27, and 35, where N4SID underestimates . The band-limited SISO approach works well when dealing with known frequency bounds, but this is never almost the case in experimental dynamics; thus, higher-order models and means of discerning between numerically spurious modes and real modes are needed. In the SIMO, MISO, and MIMO test this is tackled with stabilisation diagrams [31], where subsequent (in terms of ascending model order ) identified , , and are checked for stability to verify their physical meaningfulness. The SIMO identification obtained by iLF can identify 4 more modes (# 5, 30, 39, and 40) but lose out on mode # 11. These are expected, because as one get a more complete distribution of data along the system, more dominant (in terms of energy amplitude) modes can overshadow weaker modes (usually local or componential) close in frequency. However, a more holistic approach allows to identify modes which are not as apparent in other parts of the structure. This is a recurrent theme as more modes are identified as the test case involves more outputs and channels.
Notably, the SIMO and MISO identification is mostly unsuccessful with N4SID, as only one stable mode is identified in the MISO case, as confirmed in Figure 11(a), and none for the SIMO case. On the other hand, iLF (Figure 11(b)) can identify many new modes (8.7 % more than the SISO), in particular in the 68-84 Hz interval, and most of the SISO modes. Even more, the MIMO iLF identification can detect even more modes as it considers all 91 channels, scattered around the aircraft. This analysis identifies a total of 33 modes, five of them being at their first identification. To the Authors’ knowledge, this is the first time the MIMO modal identification of the BAE System Hawk T1A aircraft has been carried out. In general terms, the identified by different cases are coherent and show little deviation, showing confidence in the proposed iLF identification approach.
@@
Reagrding , the results for SISO identification with N4SID and iLF are mostly wrong and inaccurate. The former shows negative (Modes #11 and 17) and the latter negligible values for all modes. This is expected, as N4SID and iLF are not meant to be used in a peak-picking fashion. Also, it should be noted that the RFP implementation in [25, 30] shows negative damping for Mode #27 of the LLC-07 sensor as well. Nevertheless, the SISO values obtained from RFP compare relatively well to those obtained by the SIMO and MISO identifications, validating even more the goodness of those identifications.
@@
The computational efficiency of N4SID and iLF on the experimental SIMO and MISO data should be examined. The N4SID identification of the SIMO case took 13 h 35 min, while iLF only 13.748 s. The model order for both methods was set between 46 and 110. For the MISO identification, with the same system and same model order range, the time to identification for N4SID was almost 21 h (20 h 57 min), while iLF obtained the results in 3.098 s. The results were obtained in MATLAB 2021b running on a 14-core 2017 Apple iMac Pro with 128 GB of RAM. For reference, the same computation for the SIMO case with iLF takes 15.034 s (9.989 s for MISO) on a consumer-grade laptop (CPU 12th Gen Intel Core i5-1235U @ 1.30 GHz, 16 GB RAM, Windows 11 and MATLAB 2021b). Thus, the most complete identification of the system, the MIMO identification, is carried out solely with iLF, due to the clear computational inefficiency of N4SID. The MIMO iLF identification takes 15.084 s in the iMac Pro and 10.710 s in the consumer-grade laptop.
@@
In the stabilisation diagram in Figure 12, the stable modes can be clearly identified by the straight lines made up of the 33 stable modes identified.
@@
Having presented the identified and , the identified from the MIMO dataset need to be presented. These have been computed from all 91 accelerometer channels available; for conciseness, only the first six are shown in Figure 13 and their explanation is given in Table 4.
@@
For visualisation purposes, the landing gear channels are not portrayed and the modal displacements have been scaled with respect to a characteristic length defined depending on the accelerometer position. The wing modal displacements are scaled to the aircraft semi-span (4.7 m), and those relative to the fuselage to its maximum height (1.8 m), while the horizontal stabiliser displacements are scaled to its semi-span (2.18 m). Finally, the vertical stabiliser displacements are scaled to its span, 2.05 m. The only known size is the semi-span the others are retrieved from an existing CAD model [32].
@@
It is clear that, due to the high number of channels some artifacts are created in the modal displacement. In particular, this is seen for (Figures 13(d) and 13(f)), where the port wing tip measurement is a bit unstable. It should be noted that some sensors are not aligned to the axis system defined in Figure 9, which means that suitable rotation matrices had to be generated to reconcile the displacements with the reference system directions. In addition, the sign of some displacements had to be corrected as many are placed facing down, with respect to the positive direction of the reference system.
@@
2 | 15.43 | Dominated by horizontal stabiliser anti-symmetric (port side down) bending, and vertical stabiliser bending (port direction). Smaller contribution by wing symmetric downwards bending, |
@@
The appear more coupled than in other works, e.g. [33], due to the effect of the multi-input excitations and, in particular, of the specific setup. In fact, as remarked by the dataset authors in [26], the same driving voltage is provided to all shakers, which, in structural terms, means exciting less stiff structures (e.g. the stabilisers) with the same force as stiffer parts, e.g. the main wings. This results in an energy concentration (also due to the close proximity of the three shakers acting on the stabilisers) in a less stiff area. This is also backed up by the , as they all show a significant contribution from the stabiliser displacements. All the displacement contributions for each mode are detailed in Table 4.
All the raw data of the identified results are available in an accessible format in an open repository.

### v2

@@ 3.2 Computational performance
As already shown in [10] for the SIMO modal parameters identification, N4SID is orders of magnitude slower than LF and LSCE. LSCE is found to be the fastest method, but it is clear that it does not serve a purpose as it cannot identify all modes. However, it’s important to see that the iLF is an order of magnitude faster than the traditional implementation of LF. In addition, the conventional LF shows a slight dependency on model order, which is not the case for the iLF. The regressed line equation verifies this in Figure 3.
@@ 3.3 Investigation of noise effects
After having preliminary validated the iLF identification against the traditional LF and the benchmark methods, the next step is to assess its robustness to noise. To achieve this, the time series are corrupted with different levels, with respect to the standard deviation of the signal, of additive WGN. The selected levels are 0.1, 0.5, 1, 1.5 and 2% (note that both the input and output signals are corrupted with noise). Then, the corrupted time series are used to obtain the FRFs, which in turn are fed to iLF to obtain the modal parameters. The validation process is carried out in two steps: (i) the minimum order = 16 is considered, and (ii) the stable modes between orders 32 and 60 are examined.
@@
From Figure 4, it is clear that the first 15 modes are identified within acceptable error margins. For example, the maximum error in terms of is less than 0.5%, less than 8.5% for and the lowest MAC value of is 0.74, being 0.8 the minimum for an acceptable correlation. In addition, the first nine modes (exception made for mode #5 for the 1.5% noise level) are coherent with the analytically derived values for all noise levels. As expected, the maximum discrepancies are found for the higher modes. This is a common trend for LF-derived techniques, and it was also found in the SIMO application of LF reported in [11]. The most significant issues in modal identification precision are found for the higher modes and the modes close in frequency.
The results of the identification of stable modes from the noised signals are shown in Figure 5. Here the results are similar to that of the noiseless case, with minimal variations (less than 0.06%) in and virtually no changes in the identified (differences are all – or close to – of 0%) and (all MAC values are 1). In conclusion, the computational improvement of the iLF with respect to the conventional LF does not hamper its precision and robustness to noise. In particular, it offers the same precision as industry-standard methods, like N4SID, on the numerical cases analysed. In addition, it can identify systems that standard methods, like LSCE, struggle with - as it was for this case study. Further, experimental dataset verification and validation are carried out in the remainder of the article.
@@ 4 The BAE Systems Hawk T1A aircraft benchmark
As of the writing of this article, the experimental data from the BAE Systems Hawk T1A aircraft has been recently published by researchers from the Dynamics Research Group at the University of Sheffield. The full aircraft, a Hawk T1A formerly in service with the British Royal Air Force and now decommissioned, is depicted in Figure 6. The craft is shown at the University of Sheffield’s Structural Dynamics Laboratory for Verification and Validation (LVV) facilities (Sheffield, UK), where it was tested. Importantly, the data used in this article are firstly described in two separate documents: [29] and [30]. They focus on, respectively, the starboard wing and the whole Hawk airframe structure. Both were used here for experimental validation. These two related datasets serve as a compelling benchmark for dynamic System Identification techniques. Nevertheless, as discussed in this Section, some precautions have been considered to ensure the full comparability of data from the two experimental campaigns.
@@ 4.1 Wing experimental setup and data description
A total of 54 PCB Piezotronics accelerometers (nominal sensitivity: 10 mVg-1) are mounted on the lower and upper surface of the starboard wing, as shown in Figure 7. They are arranged in a grid of three and two lines, respectively, running parallel to the leading edge. Extra sensors are placed at the root boundary condition locations and by the undercarriage door. One Tira TV 51140-MOSP modal shaker, directly attached to a PCB Piezotronics 208C02 force transducer with a sensitivity of 11.11 mVN-1, applies the input driving force on the wing lower surface. Specifically, the acquisitions corresponding to burst random excitation – white noise forcing input, with amplitude ramping at the beginning and end of the burst – with 0.4 V LMS input voltage are considered in this study. A Siemens PLM LMS SCADAS unit was used to generate the input signal, and record the acceleration responses.
As can be inferred, this dataset only includes one input source. Therefore, it was used to obtain the SISO readings to compare the SISO and SIMO capabilities of the iLF while getting baseline values for system modes.
4.2 Whole airframe setup and data description
@@
To ensure comparability with the wing dataset, the recordings corresponding to a WGN input with the same input voltage (0.4 V - before amplification) have been used. It should be noted, however, that MIMO tests, by definition, result in an overall higher excitation energy. Here, unique drive signals for each of the five shakers were generated by summing sine waves with random phases, followed by iterative adjustments to meet a defined tolerance of 4% between measured and target power spectral densities. Once aligned, spectral gain masks were stored and used to generate the drive signals, allowing for variable test periods through frequency domain resampling by linear interpolation. In [30], no information seems to be explicitly reported on the data collection system and the software used. Nevertheless, these can be inferred to be National Instruments hardware with LabView software, since a similar control algorithm implementation is used in [29] for the closed-cloop wing tests. In this case, five modal shakers are attached to the starboard wing (i.e. the one already discussed above), the port wing, plus the starboard horizontal stabiliser, port horizontal stabiliser, and vertical stabiliser. The exact locations are marked in orange in Figures 8 and 9.
@@
The two sides of the horizontal stabiliser are populated with two rows of five uniaxial sensors at both the rear and front side, totalling 20 per side (located exclusively on the upper; see Figure 9(a));
@@
Similarly, the vertical stabiliser sensors are placed only on the starboard side, again with 10 sensors equally distributed along the leading and trailing edge (Figure 9(b));
@@
Another set of four high-sensitivity (100 mVg-1) uniaxial PCB Piezotronics accelerometers are placed on the fuselage midpoint on the upper, lower, port, and starboard sides (F_MLU, F_MLL, F_MLP, and F_MLS in Figure Figure 10(a));
@@
The 86th sensor, a triaxial accelerometer laid on the ground as a reference, is obviously not used here for SI. Apart from this, all the other output channels are selected for MIMO System Identification, in conjunction with the five input channels. Conversely, the fibre-Bragg grating strain gauges - described in [32] - are not considered here.
@@ 4.2.1 Pre-processed FRFs
estimators are indeed appreciated for their capability to reduce output measurement noise, which more significantly affects the valleys of FRFs with signal leakage. However, it is well-known that, while the FRF estimators allow for averaging over multiple acquisitions and provide more accurate estimates in absolute value, this is achieved at the cost of losing the phase information in the auto-spectra. Nevertheless, this information is retained in the cross-spectra (see e.g. [33] Chapter 3.13).
@@ 4.2.2 Further considerations
Before continuing, there are two key aspects worth mentioning. The first one is that the sensor layout for the starboard wing used in [32] is similar but not identical to the one of [31], as e.g. only two lines of sensors were used on the upper side (one on the lower side). As it is better discussed in the Section 4, the nearest output channels were selected to compare SISO, MISO, and MIMO tests between these two datasets.
The second aspect is that for the experiments on the whole airframe with WGN input, the excitation bandwidth was 5-256 Hz and the sampling rate was 2048 Hz. On the other hand, the first experimental campaign used a sampling frequency of 512 Hz, but it was bandlimited between 5-256 Hz. Therefore, all analyses have been here limited to Hz.
@@ 4.3 Modal parameters estimation
Previous MIMO modal testing work on the aircraft does not explicitly report any results to be used as an external benchmark [30]. Nevertheless, data is available from the SIMO modal testing of the aircraft starboard wing [29]. However, the results presented here are only for the SISO identification via RFP from the FRF for channel LTC-05 (Figure 11(a)), with no further numerical values included. It is clear that, for getting an idea of where (in frequency) the modes of the full aircraft are, this approach is limited by the data considered. In addition, it should be noted that the second to last mode (mode #22 according to [29]) is very likely a repetition of the previous mode (potentially a typo). Nevertheless, it is possible to retrieve its actual value, as the authors shared the code used to obtain those results in [34].
In [29], ranges for the natural frequencies of the starboard wing are supplied to be used to identify the modes via RFP; hence, the identification is carried out in a peak-picking fashion rather than using higher-order models. Thus, the idea is to compare this SISO result with a MISO result from the same, or close, sensor. To do this, the sensors position graphs of the starboard wing supplied in [29] (SIMO) and [30] (MIMO) are superimposed in Figure 11(a). There, it is shown that the shaker is located in (almost) the same position in the two tests; however, the sensor row of LTC-05 is not present in the MIMO sensor layout. Hence, another sensor, with a counterpart in the MIMO test, is identified. In the SIMO test, this sensor is known as LLC-07, while in the MIMO test, it is identified as SW_LC1.
@@
The expectation is that N4SID and iLF will not perform as well as RFP in this task as they are more holistic methods implemented for more typical scenarios with multiple outputs and inputs rather than being used in a peak-picking fashion. Then, the SIMO and MISO analyses outlined above follow.
The identified and , for the different cases and sensors, are shown, respectively, in Tables 3 and 4. The results marked with (*) are those erroneously reported in [29] and retrieved with the authors-provided [34] RFP implementation. Please note that only frequencies 5-160 Hz are inspected in all cases for coherence with [29].
@@
The results presented in [29] for the band-limited SISO identification via RFP of the LTC-05 sensor of the starboard wing dataset show that 23 modes are found for the wing. Of course, these modes are representative of a local state, and more modes are found when the full aircraft is tested. After all the SISO, SIMO, MISO, and MIMO identifications, a total of 42 modes are identified in at least one of the cases. Hence, from now on, these modes are referred to with the numbering shown in Tables 3 and 4.
In terms of for the SISO identifications, N4SID and iLF always match the frequencies identified by RFP in LTC-05 and LLC-07, except for modes # 20, 27, and 35, where N4SID underestimates . The band-limited SISO approach works well when dealing with known frequency bounds, but this is never almost the case in experimental dynamics; thus, higher-order models and means of discerning between numerically spurious modes and real modes are needed. In the SIMO, MISO, and MIMO tests, this is tackled with stabilisation diagrams [35], where subsequent (in terms of ascending model order ) identified , , and are checked for stability to verify their physical meaningfulness. The SIMO identification obtained by iLF can identify 4 more modes (# 5, 30, 39, and 40) but lose out on mode # 11. These are expected because as one gets a more complete distribution of data along the system, more dominant (in terms of energy amplitude) modes can overshadow weaker modes (usually local or componential) close in frequency. However, a more holistic approach allows to identify modes which are not as apparent in other parts of the structure. This is a recurrent theme as more modes are identified as the test case involves more outputs and channels.
Notably, the SIMO and MISO identification is mostly unsuccessful with N4SID, as only one stable mode is identified in the MISO case, as confirmed in Figure 12(a), and none for the SIMO case. On the other hand, iLF (Figure 12(b)) can identify many new modes (8.7 % more than the SISO), in particular in the 68-84 Hz interval, as well as most of the SISO modes. Even more, the MIMO iLF identification can detect even more modes as it considers all 91 channels located all around the aircraft. This analysis identifies a total of 33 modes, five of them being at their first identification. To the authors’ knowledge, this is the first time the MIMO modal identification of the BAE System Hawk T1A aircraft has been carried out. In general terms, the identified by different cases are coherent and show little deviation, showing confidence in the proposed iLF identification approach.
@@
Regarding , the results for SISO identification with N4SID and iLF are mostly wrong and inaccurate. The former shows negative (Modes #11 and 17), and the latter negligible values for all modes. This is expected, as N4SID and iLF are not meant to be used in a peak-picking fashion. Also, it should be noted that the RFP implementation in [29, 34] shows negative damping for Mode #27 of the LLC-07 sensor as well. Nevertheless, the SISO values obtained from RFP compare relatively well to those obtained by the SIMO and MISO identifications, validating the goodness of those identifications even more.
@@
Next, the computational efficiency of N4SID and iLF on the experimental SIMO and MISO data should be examined. The N4SID identification of the SIMO case took 13 h 35 min, while iLF only 13.748 s (a reduction of circa 99.97% of the elapsed time). The model order for both methods was set between 46 and 110. For the MISO identification, with the same system and same model order range, the time to identification for N4SID was almost 21 h (20 h 57 min), while iLF obtained the results in 3.098 s (circa 99.99% reduction). The results were obtained in MATLAB 2021b running on a 14-core 2017 Apple iMac Pro with 128 GB of RAM. For reference, the same computation for the SIMO case with iLF takes 15.034 s (9.989 s for MISO) on a consumer-grade laptop (CPU 12th Gen Intel Core i5-1235U @ 1.30 GHz, 16 GB RAM, Windows 11 and MATLAB 2021b). Thus, the most complete identification of the system, the MIMO identification, is carried out solely with iLF due to the evident less computational efficiency of N4SID. The MIMO iLF identification takes 15.084 s on the iMac Pro and 10.710 s on the consumer-grade laptop.
@@
Figure 13 show the stabilisation diagram for the MIMO identification via iLF. This clearly shows the straight lines made up of the 33 stable modes.
@@
Having presented the identified and , the identified from the MIMO dataset need to be discussed. These have been computed from all 91 accelerometer channels available; for conciseness, only the first six are shown in Figure 14, and their explanation is given in Table 5.
@@
For visualisation purposes only, the landing gear channels are not portrayed. The modal displacements have been scaled with respect to a characteristic length defined depending on the accelerometer position. The wing modal displacements are scaled to the aircraft semi-span (4.7 m), those relative to the fuselage to its maximum height (1.8 m), and those of the horizontal stabiliser displacements to its semi-span (2.18 m). Finally, the vertical stabiliser displacements are scaled to its span, 2.05 m. The only known size is the semi-span the others are retrieved from an existing CAD model [36].
@@
It is clear that, due to the high number of channels some artifacts are created in the modal displacement. In particular, this is seen for (Figures 14(d) and 14(f)), where the port wing tip measurement is quite unstable. It should be noted that some sensors are not aligned to the axis system defined in Figure 10, which means that suitable rotation matrices had to be generated to reconcile the displacements with the reference system directions. In addition, the signs of some displacements had to be corrected as many were likely placed facing down, with respect to the positive direction of the reference system.
@@
2 | 15.43 | Dominated by horizontal stabiliser anti-symmetric (port side down) bending, with vertical stabiliser bending (port direction). Smaller contribution by wing symmetric downwards bending, |
@@
The appear more coupled than in other works, e.g. [37], due to the effect of the multi-input excitations and, in particular, of the specific setup. In fact, as remarked by the dataset authors in [30], the same driving voltage is provided to all shakers, which, in structural terms, means exciting less stiff structures (e.g. the stabilisers) with the same force as stiffer parts, e.g. the main wings. This results in an energy concentration (also due to the close proximity of the three shakers acting on the stabilisers) in a less stiff area. This is also backed up by the , as they all show a significant contribution from the stabiliser displacements. All the displacement contributions for each mode are detailed in Table 5. All the raw data of the identified results are available in an accessible format in an open repository.